OpenStudy (idealist10):

Find all solutions of xy'=2-x+(2x-2)y-xy^2.

OpenStudy (idealist10):

xy'=2-x+2xy-2y-xy^2 xy'=-(xy^2-2xy+x)+2(1-y) xy'=-x(y-1)^2-2(y-1) y'=-(y-1)^2-2(y-1)/x u=y-1 u'=y' u'=-u^2-2u/x w=ux What should I do now? w=ux is the substitution that should give me a separable DE.

OpenStudy (idealist10):

@hartnn @ganeshie8

hartnn (hartnn):

w = ux or w= u/x ?

OpenStudy (idealist10):

I need to find w' but what's w'? I know that it's the derivative of w and I need to use the product rule but what's w'?

OpenStudy (idealist10):

w=ux

hartnn (hartnn):

w= ux product rule w' = u x' + u'x w' = u +u'x

OpenStudy (idealist10):

How did you get w'=u+u'x from w'=ux'+u'x?

ganeshie8 (ganeshie8):

y = f(x) u = y-1 = f(x) - 1 w = ux = function of x too

hartnn (hartnn):

w=ux product rule \((fg)' = f'g +fg'\) \((ux)' = u'x+ux'\) x'=1

ganeshie8 (ganeshie8):

the independent variable is x in all your substitutions

ganeshie8 (ganeshie8):

you're assuming implicitly : \(y' = \dfrac{dy}{d\color{Red}{x}}\) \(u' = \dfrac{du}{d\color{Red}{x}}\) \(w' = \dfrac{dw}{d\color{Red}{x}}\)

OpenStudy (idealist10):

Now I understand how w'=u+u'x but what do I do next? I need to get a separable DE. I'm so sorry for the slowness of my internet, @hartnn @ganeshie8

ganeshie8 (ganeshie8):

u'=-u^2-2u/x i would substitute u = wx

ganeshie8 (ganeshie8):

then you get : u' = w'x + w

hartnn (hartnn):

w=u/x work ? in my attempt, it didn't...

ganeshie8 (ganeshie8):

Oh the given equation is not homogeneous to start with

ganeshie8 (ganeshie8):

it wont work yeah

hartnn (hartnn):

right, we need to identify the form of DE ..

hartnn (hartnn):

first order non-linear

OpenStudy (idealist10):

So does the substitution u=wx and u'=w+w'x works?

hartnn (hartnn):

w=ux was given as a part of hint or something or its your try ?

OpenStudy (idealist10):

Part of hint.

hartnn (hartnn):

u' = (xw' -w)/x^2

hartnn (hartnn):

= -w^2 /x^2 -2w/x^2 xw' -w = -w^2 -2w too much algebra

hartnn (hartnn):

but finally , magically, its in the separable form!

hartnn (hartnn):

so yes w=ux worked

hartnn (hartnn):

could have never guessed it without the hint :P

OpenStudy (idealist10):

So I got xw'=-w(w+1)

hartnn (hartnn):

yep

hartnn (hartnn):

x dw/dx = -w (w+1) dx/x = - dw/ w(w+1) quite simple

hartnn (hartnn):

did u get this : u' = (xw' -w)/x^2

OpenStudy (idealist10):

Yes. :)

ganeshie8 (ganeshie8):

nice :)

hartnn (hartnn):

then you should not have any problems solving it further :)

OpenStudy (idealist10):

Did you know that I've struggled with this problem since I got u'=(xw'-wx')/x^2 instead of u'=(xw'-w)/x^2? HAHAHA, I'm so stupid. Thank you so much, @hartnn !!

hartnn (hartnn):

lol happens :)